how many gallons of 70% anifreeze solution must be mixed with 70 gallons of 15% antifreeze to get a mixture that is 60% antifreeze?
Answer by htmentor(1343)

how many gallons of 70% anifreeze solution must be mixed with 70 gallons of 15% antifreeze to get a mixture that is 60% antifreeze?
==========================================================
Let x = the number of gallons of 70% antifreeze solution
Then the number of gallons of antifreeze making up the mixture will be:
0.7x + 0.15*70 = 0.6*(x+70)
0.7x + 10.5 = 0.6x + 42
Solve for x:
0.1x = 31.5
x = 315
So, we need 315 gallons of 70% antifreeze solution